OpenStudy (anonymous):

What is the exact value of the expression . 6 sqrt y^8-sqrt 9y^4+ 3 sqrt y^8 - sqrt y^4 Simplify if possible. Please with steps

OpenStudy (anonymous):

\[6\sqrt{y^8}=6y^4\] \[-\sqrt{9y^4}=-3y^2\] \[3\sqrt{y^8}=3y^4\] \[-\sqrt{y^4}=-y^2\] combine like terms for the last step